OpenStudy (anonymous):

What is the simplified form of 1/10xy multiplied by 5x^2/7y

OpenStudy (maheshmeghwal9):

\[(1/10xy)*(5x^2/7y)=(x/2)*(1/7y)=(x/14y).\]

OpenStudy (maheshmeghwal9):

sorry! (x/2y)*(1/7y)=(x/14y^2)

OpenStudy (anonymous):

thanks

OpenStudy (maheshmeghwal9):

wc!